blabla:

A quarterback looks downfield at a receiver who is 14 yards away. The receiver is covered, so the quarterback turns 30° to his right and sees another receiver who is 19 yards away. To the nearest tenth of a yard, how far apart are the two receivers?

TheSmartOne:

Hey there! Welcome to QuestionCove! Do you know the law of cosines? That's what we'll have to use to solve your question If you don't remember it, you can quickly brush up on it here before we plug and chug into the formula https://www.onlinemathlearning.com/law-of-cosines.html

TheSmartOne:

\(\ a^2 = b^2 + c^2 - 2bc\cos A\)

TheSmartOne:

As long as we know the value of two sides and the angle in between them, we can use the law of cosines to find the third side

TheSmartOne:

This should help you visualize it better https://www.onlinemathlearning.com/image-files/xlaw-of-cosines.png.pagespeed.ic.xmYFXUz4Xc.png